OpenStudy (anonymous):

The height of a triangle is twice the base. The area is 121 square inches. Find the base. @ganeshie8

OpenStudy (anonymous):

Same routine as the last question you showed

OpenStudy (anonymous):

Area of triangle = 1/2bh Height = 2b

ganeshie8 (ganeshie8):

plugin the value of height in area equation

ganeshie8 (ganeshie8):

121 = 1/2bh 121 = 1/2b(2b) solve b

OpenStudy (anonymous):

uhhh, 121=1/4? I'm lost

ganeshie8 (ganeshie8):

\(121 = \dfrac{1}{2} b(2b)\)

ganeshie8 (ganeshie8):

\(121 = \dfrac{1}{\not 2} b(\not 2b)\)

ganeshie8 (ganeshie8):

\(121 = b^2\)

OpenStudy (anonymous):

Oooooh!

ganeshie8 (ganeshie8):

\(? = b\)

OpenStudy (anonymous):

So, we have to get rid of the square?

ganeshie8 (ganeshie8):

too easy for u cookie

ganeshie8 (ganeshie8):

yes just take square root both sides

OpenStudy (anonymous):

11!!!! lol

OpenStudy (anonymous):

It's easy & I still don't understand lol

ganeshie8 (ganeshie8):

its okay dont wry, someday everything wil make sense lol :P

OpenStudy (anonymous):

If it hasn't by now, it never will lol

OpenStudy (anonymous):

YOOOO not even, I'm terrible at math, but look at my SS now :3

OpenStudy (anonymous):

lol mines, not going past 60 lol Mines only go up from asking questions & getting fans lol

ganeshie8 (ganeshie8):

hahah cookies need love...

ganeshie8 (ganeshie8):

learn it wid love, u wil enjoy and everything wil make sense... :) thats all i can say for now lol

OpenStudy (anonymous):

I HATE math lol It even says that in my bio lol

OpenStudy (anonymous):

Believe in yourself cookii, I'm actually bad at math, but when I see ganeshie8 it just comes to me, he has that kind of a impact on me :). But, honestly it takes lots of practice, we were all in the position where you are, actually I was way worse.

OpenStudy (anonymous):

@ganeshie8 is sooooo much help !! lol We're going to be bestfriends lol

ganeshie8 (ganeshie8):

woah ty for the kind words lol xD cookie why are u doing math everyday if you hate math ?

OpenStudy (anonymous):

College :/ It's all online & it's the only class I ever have homework in lol

ganeshie8 (ganeshie8):

lol ikr... since you're forced to work on math everyday... isnt it a clever idea to approach it wid some affection...giving it hugs and if psble kisses each time u open the math homework.. instead of cursing it and treating it like s@#t ? :)

OpenStudy (anonymous):

lmaoo F U C K math lol

OpenStudy (anonymous):

We all hate math sometimes, it bullies me all the time when I'm wrong, but that makes me want to just find the right answer even more >:)

OpenStudy (anonymous):

Makes me want to quit & watch Pysch, but since I finished all the episodes & the show got canceled, I have to -_-